20,000 hrs                        

30,000 hrs                      

40,000 hrs                                               

variable costs                            

?

?

?

Variable cost per hr

?

?

$3.00                                                          

fixed costs

?

$200,000

?

fixed cost per hr

?

?

?

total costs

?

?

?

total cost per hr                          

?

?

?

Fill in the missing numbers.

What would be the variable costs and fixed costs at 25,000 hours of activity?
Answer #1

Fixed cost in totality remains same but would decrease at per unit level if production increase on the other hand variable cost increase and decrease in totality but remains same at unit level.

Solution

20,000 hrs                        

30,000 hrs                      

40,000 hrs                                               

variable costs                            

$      60,000.00

$     90,000.00

$ 1,20,000.00

Variable cost per hr

$                 3.00

$                3.00

$                3.00

fixed costs

$ 2,00,000.00

$ 2,00,000.00

$ 2,00,000.00

fixed cost per hr

$              10.00

$                6.67

$                5.00

total costs

$ 2,60,000.00

$ 2,90,000.00

$ 3,20,000.00

total cost per hr                          

$              13.00

$                9.67

$                8.00

Working

20,000 hrs                        

30,000 hrs                      

40,000 hrs                                               

variable costs                            

=3*20000

=3*30000

=3*40000

Variable cost per hr

3

3

3

fixed costs

200000

200000

200000

fixed cost per hr

=200000/20000

=200000/30000

=200000/40000

total costs

=200000+60000

=200000+90000

=200000+120000

total cost per hr                          

=260000/20000

=290000/30000

=320000/40000